# Search for Light Dark Matter with 259-day data in PandaX-4T Minzhen Zhang

### v1

@@
Search for Light Dark Matter with 259-day data in PandaX-4T
@@ Abstract
We establish the most stringent constraints on spin-independent dark matter-nucleon interactions within a mass range of 2.5 to 5.0 GeV/, spin-dependent neutron-only interactions within 1.0 to 5.6 GeV/, and spin-dependent proton-only interactions within 1.0 to 4.1 GeV/.
Additionally, our results improve the upper limits on the dark matter-electron scattering cross-section by a factor of 1.5 and 9.3 for heavy and light mediator scenarios respectively within 50 MeV/ to 10 GeV/, compared with previous best results.
Large amount of astrophysical observations have supported the common existence of dark matter (DM) in our Universe [Zwicky_1933gu, Rubin_1970zza, Umetsu_2011, Clowe_2006, Planck2016], but the nature of DM is still unknown.
@@
The most probable mass range of such cold DM particles is between about several GeV/ to approximately 10 TeV.
Decades-long searches for WIMPs have not yielded conclusive detections yet. However, they led to the development of various ultra-sensitive detection techniques [dm_direct_detection_review].
@@
It promotes studies of the light DM (LDM) particles which have masses approximately from keV/ to sub-GeV/. There are many LDM candidates, including sterile neutrinos [Boyarsky:2009ix], axions and axion-like-particles [Marsh:2015xka], hidden sector particles [Feng:2011ik], dark photons [Feldstein:2013kka], superWIMPs [Feng:2003uy] and DM from primordial black hole evaporation [Masina:2020xhk], etc. LDM, produced via freeze-in, freeze-out, or non-thermal mechanisms, typically interacts weakly with the Standard Model particles, addressing small-scale structure issues and offering diverse detection signatures in astrophysical and cosmological observations [Arcadi:2017kky].
@@
The PandaX-4T [PandaX-4T:2021bab, PandaX:2024qfu], located at the China Jinping Underground Laboratory (CJPL) [kang2010status, li2015second], is a deep-underground multi-tonne liquid xenon detector designed to search for ultra-rare events from dark matter particles, Majorana neutrinos, and astrophysical neutrinos.
It utilizes a dual-phase cylindrical time projection chamber (TPC) with a diameter of 1.2 meters and a height of 1.2 meters.
The detector uses approximately 5.6 tonnes of liquid xenon as the target material, with around 3.7 tonnes within the sensitive volume.
@@
(Top) The differential rates (in unit of events per tonne per year per keV) of the DM scatterings off the target.
@@
The right panel shows the spectra for LDM-electron scatterings with LDM masses of 50 MeV/c2 (red) and 0.1 GeV (blue), assuming an LDM-electron cross section of 10-37 cm2.
@@
(Middle) All components of the selection efficiency for the paired (solid) and US2 (dashed) data: green=data quality selection, blue=signal reconstruction, cyan=ROI, magenta=BDT, black=total.
@@
(Bottom) The charge (black) yields as a function of NR and ER energies are shown in the left and right panel, respectively.
@@
However, there may also be spin-dependent (SD) contribution to the DM-nucleus cross section, like axial-vector mediated interaction, etc.
Only nuclear isotopes with non-zero spin can participate in the SD scattering processes.
@@
where , , and are the local DM density, the mass of the DM particle, and the reduced mass between DM and the target nucleus.
@@
For simplicity, this work focuses on the proton-only and neutron-only cases to compare with constraints presented in the literature. Given the fact that xenon nucleus has an even number of protons, the spin of xenon nucleus is dominated by the unpaired neutron due to the spin pairing effects and the event rate of the neutron-only interaction is significantly larger than the proton-only case even with the same DM-nucleon cross section.
@@
The differential rate for the LDM-electron scattering can be expressed as integral over (n,l) shell contributions [essig2017new]:
@@
is the inverse mean speed, is the reduced LDM-electron mass, and is the momentum transfer dependent LDM-electron form factor.
@@
The differential rates of the LDM-nucleus and LDM-electron scatterings calculated based on Eqs. 1 and 2 are shown in the top panels of Fig. 1.
@@
This combined approach enables sensitivity across a wide energy range: approximately 0.33–3.0 keV for nuclear recoils (NRs) and 0.04–0.39 keV for electronic recoils (ERs).
@@
For the paired data, we require the signal to be detected by 2 or 3 PMTs, while the US2 data must have the number of electrons for an to be between 4 to 8 detected electrons and no signal detected by 2 or more PMTs in the event window.
The total efficiencies, including contributions from the quality selections, the BDT cuts, the region-of-interest (ROI) cuts, and the signal reconstruction, as a function of NR and ER energies, are presented in the middle panels of Fig.1.
@@
Top panel: Data and fitted (for a LDM of 4 GeV/c2) spectra in number of electrons, Run0 and Run1 combined.
@@
Middle and bottom panels: The US2 data(black dots) distribution of the width versus charge. Along with expected 1 contours of 4 GeV/c2 DM-nucleon spin-independent interaction(red dashed line), 2 GeV/c2 DM-electron interaction via heavy mediator(blue dashed line) and light mediator(black dashed line), 8B solar neutrino(magenta shadow), MD background(yellow shadow) and cathode background(green shadow).
@@ 2 GeV/c2
Table 1: The expected and best-fit events number of dark matter, 8B solar neutrino, AC background, Cathode background and MD background in paired and US2 channels, assuming 4 GeV/c2 DM-nucleon spin-independent interaction, 2 GeV/c2 LDM-electron interaction via heavy mediator and light mediator.
@@
In the paired dataset, the dominant background arises from accidental pileup (AC) events, whereas the US2 dataset is primarily affected by three components: the cathode-originated background, the pileup of delayed electrons (MD background), and the coherent elastic neutrino-nucleus scattering induced by solar 8B neutrinos (8B CENS).
@@
The models have been tested with a side band before unblinding as in Ref. [PandaX:2022xqx] and Ref. [PandaX:2022aac].
Systematic uncertainties are assigned to the selection acceptance, BDT efficiency, MD background rate, cathode background rate, and signal response model, following the same methodology of Ref. [bo2024first].
@@
The signal response models for LDM and 8B CENS interactions in both the paired and US2 data are constructed using the PandaX-4T signal response framework (bottom panels of Fig. 1) [luo2024signal].
@@
For comparison, we also interpret the results using the charge yields from NEST [szydagis2022review], leading to slightly more aggressive results.
Both the NR and ER signal model uncertainties are considered.
In the paired (US2) dataset, we observe 1 (158) and 2 (174) events in Run0 and Run1, respectively.
@@
The joint (, ) distributions of the MD and cathode backgrounds are derived from controlled samples in the experimental data, while those for DM and 8B CENS signals are generated by the signal response model, assuming a uniform spatial distribution within the FV.
Fig. 2 displays the 1 confidence contours for both background components and DM signals with masses of 4 GeV/c2 (SI DM-n) and 2 GeV/c2 (LDM-e), alongside the observed US2 event distributions from Run0 and Run1.
@@
Since there are almost no expected LDM-electron scattering events in paired datasets, this channel is not included in statistical analysis of LDM-electron interaction.
@@
The 90% C.L. exclusion limit(blue line) with 1 sensitivity(green band) on spin-independent DM-nucleon cross section(left pannel), neutron-only(middle pannel) and proton-only(right pannel) spin-dependent DM-nucleon cross section, together with results from other work [PandaX:2022xqx, PandaX:2022aac, XENON:2024hup, XENON:2019gfn, DarkSide:2018kuk, LUX:2016ggv, PandaX-II:2018woa, PICO:2019vsc].
@@
The 90% C.L. exclusion limit(blue line) with 1 sensitivity(green band) on DM-electron cross section via heavy(top pannel) and light mediator(bottom pannel) using constant-W model. The aggressive results using NEST light and charge yields are also overlaid(blue dashed line), as well as results of other experiments [PandaX:2022xqx, XENON:2019gfn, DarkSide:2022knj, adari2025first, aggarwal2025probing], and cosmological predictions from DM vector-portal freeze-in and freeze-out mechanisms [Essig:2015cda].
@@
Figure 3 presents the 90% confidence level (C.L.) exclusion limits on the DM-nucleon scattering cross sections.
The results establish the most stringent limits for SI, SD (neutron-only), and SD (proton-only) DM-nucleon scatterings in the mass range of approximately [2.5, 5.0], [1.0, 5.6], and [1.0, 4.1] GeV/c2, respectively.
The corresponding exclusion limits at 3 GeV/c2 are , , and cm2.
@@
The SI limits exhibit stronger upward fluctuations above 3.5 GeV/c2 because the harder energy spectrum populates the region (5 8 electrons) where we observe a slight event excess.
This trend also holds for SD limits due to their typically hard spectra.
Below 3.5 GeV/c2, SI limits show no significant upward fluctuations as the spectra become dominated by MD background.
The LDM-electron cross section exclusion limits are improved by approximately 1.5 times for the heavy-mediator case, and 9.3 times for the light-mediator case, in mass range of [50, 10000] MeV/c2.
The improvement in light-mediator case is mainly because of different treatments in signal model 111
In previous work [PandaX:2022xqx], the mean total number of created photons and electrons was conservatively set to a floored integer.
This treatment has a significant impact on the spectrum near our ROI threshold, particularly affecting searches for LDM–electron interactions mediated by a light mediator.
The limits for LDM with heavy mediators exhibit stronger upward fluctuations than those with light mediators, as heavy-mediator spectra predominantly populate the excess region (5 8 electrons), while light-mediator spectra primarily occupy the MD region (4 5 electrons).
In addition, due to the expected events gathering in the low energy region where MD is dominant, greatly reduced MD uncertainty(from 130% [PandaX:2022xqx] to 13% for Run0 and 16% for Run1) also improves the limits by around 2 times in light-mediator case.
The limit now has excluded portion of the cosmologically allowed parameter space from the cosmic microwave background measurements [Essig:2015cda], in the mass range from about 50 MeV/c2 to 10 GeV/c2, which is shown in Fig. 4.
@@
Our analysis establishes new exclusion limits at 90% C.L. for: spin-independent DM-nucleon scattering in the mass range [2.5, 5.0] GeV/c2, spin-dependent DM-nucleon scattering ([1.0, 5.6] GeV/c2 for neutron-only coupling, [1.0, 4.1] GeV/c2 for proton-only coupling), and LDM-electron scattering (from 100 MeV/c2 to 10 GeV/c2).
@@
For the LDM model, our results constrain previously unexplored parameter space given by the cosmological predictions [Essig:2015cda].
@@
This project is supported in part by grants from National Key R&D Program of China (Nos. 2023YFA1606200, 2023YFA1606201), National Science Foundation of China (Nos. 12090060, 12090061, 12090063, 12325505, 12222505, 12275267, U23B2070), and by Office of Science and Technology, Shanghai Municipal Government (grant Nos. 22JC1410100, 21TQ1400218), Sichuan Province Innovative Talent Funding Project for Postdoctoral Fellows (No. BX202322), Sichuan Provincial Natural Science Foundation (No. 2024NSFSC1374). We thank for the support by the Fundamental Research Funds for the Central Universities. We also thank the sponsorship from the Chinese Academy of Sciences Center for Excellence in Particle Physics (CCEPP), Hongwen Foundation in Hong Kong, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund. Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.

### v3

# Search for Light Dark Matter with 259 Days of Data in PandaX-4T Minzhen Zhang

@@
Search for Light Dark Matter with 259 Days of Data in PandaX-4T
@@ Abstract
We establish the most stringent constraints on spin-independent dark matter-nucleon interactions within a mass range of 2.5-5.0 GeV/, spin-dependent neutron-only interactions within 2.0-4.6 GeV/, and spin-dependent proton-only interactions within 2.0-3.6 GeV/.
Their corresponding limits at 3 GeV/ are , , and cm2, respectively.
Additionally, our results improve the upper limits on the dark matter-electron scattering cross section by a factor of 1.5 and 9.3 for heavy and light mediator scenarios, respectively, within 50 MeV/ to 10 GeV/, compared with previous best results.
The improved limits reach 1.510-41 and 5.610-37cm2 at 200 MeV/ for the heavy and light mediators, respectively.
A Large number of astrophysical observations have supported the common existence of dark matter (DM) in our Universe [Zwicky_1933gu, Rubin_1970zza, Umetsu_2011, Clowe_2006, Planck2016], but the nature of DM is still unknown.
@@
The most probable mass range of such cold DM particles is between about several GeV/ and approximately 10 TeV.
Decades-long searches for WIMPs have not yielded conclusive detections yet. However, they led to the development of various ultrasensitive detection techniques [dm_direct_detection_review].
@@
They promote studies of the light DM (LDM) particles which have masses approximately from keV/ to sub-GeV/. There are many LDM candidates, including sterile neutrinos [Boyarsky:2009ix], axions and axionlike particles [Marsh:2015xka], hidden sector particles [Feng:2011ik], dark photons [Feldstein:2013kka], superWIMPs [Feng:2003uy], and DM from primordial black hole evaporation [Masina:2020xhk], etc. LDM, produced via freeze-in, freeze-out, or nonthermal mechanisms, typically interacts weakly with the standard model particles, addressing small-scale structure issues and offering diverse detection signatures in astrophysical and cosmological observations [Arcadi:2017kky].
@@
The PandaX-4T [PandaX-4T:2021bab, bo2025dark], located at the China Jinping Underground Laboratory (CJPL) [kang2010status, li2015second], is a deep-underground multitonne liquid xenon detector designed to search for ultrarare events from dark matter particles, Majorana neutrinos, and astrophysical neutrinos.
It utilizes a dual-phase cylindrical time projection chamber (TPC) with a diameter of 1.2 m and a height of 1.2 m.
The detector uses approximately 5.6 metric tons of liquid xenon as the target material, with around 3.7 metric tons within the sensitive volume.
@@
Top: the differential rates (in unit of events per metric ton per year per keV) of the DM scatterings off the target.
@@
The right panel shows the spectra for LDM-electron scatterings with LDM masses of 50 MeV/ (red) and 0.1 GeV (blue), assuming an LDM-electron cross section of 10-37 cm2.
@@
All spectra are without efficiency applied.
Middle: all components of the selection efficiency for the paired (solid) and US2 (dashed) data: green, data quality selection; blue, signal reconstruction; cyan, ROI; magenta, BDT; black, total.
@@
The efficiencies as a function of NR and ER energies are estimated based on the nominal charge yields shown in the bottom panels.
Bottom: the charge (black) yields as a function of NR and ER energies are shown in the left and right panel, respectively.
@@
No significant deviation from the nominal models is observed.
@@
However, there may also be a spin-dependent (SD) contribution to the DM-nucleus cross section, like axial-vector mediated interaction, etc.
Only nuclear isotopes with nonzero spin can participate in the SD scattering processes.
@@
where , , and are the local DM density, the mass of the DM particle, and the reduced mass between DM and the target nucleus, respectively.
@@
For simplicity, this work focuses on the proton-only and neutron-only cases to compare with constraints presented in the literature. Given the fact that the xenon nucleus has an even number of protons, the spin of a xenon nucleus is dominated by the unpaired neutron due to the spin pairing effects and the event rate of the neutron-only interaction is significantly larger than the proton-only case even with the same DM-nucleon cross section.
@@
The differential rate for the LDM-electron scattering can be expressed as integral over shell contributions [essig2017new]:
@@
is the inverse mean speed, is the reduced LDM-electron mass, and is the momentum transfer-dependent LDM-electron form factor.
@@
The differential rates of the LDM-nucleus and LDM-electron scatterings calculated based on Eqs. 1 and 2 are shown in the top panels in Fig. 1.
@@
This combined approach enables sensitivity across a wide energy range: approximately 0.33–3.0 keV for NRs and 0.04–0.39 keV for ERs.
@@
For the paired data, we require the signal to be detected by two or three PMTs, while the US2 data must have the number of electrons for an to be between four and eight detected electrons and no signal detected by two or more PMTs in the event window.
The total efficiencies, including contributions from the quality selections, the BDT cuts, the region-of-interest (ROI) cuts, and the signal reconstruction, as a function of NR and ER energies, are presented in the middle panels in Fig.1, and can also be found in the Supplementary Material [SupplementaryMaterial].
@@
The efficiencies of paired data for ER are not given, since there are almost no expected LDM-electron scattering events in paired datasets due to the much lower light-to-charge ratio for the ER signals.
This channel is not included in statistical analysis of LDM-electron interaction.
@@
Top panel: US2 data and fitted (for a LDM of 4 GeV/) spectra in number of electrons, run0 and run1 combined.
@@
Middle and bottom panels: the US2 data(black dots) distribution of the width versus charge. Along with expected 1 contours of 4 GeV/ DM-nucleon spin-independent interaction(red dashed line), 2 GeV/ DM-electron interaction via heavy mediator(blue dashed line), and light mediator(black dashed line), 8B solar neutrino(magenta shadow), MD background(yellow shadow), and cathode background(green shadow).
@@ 2 GeV/
Table 1: The expected and best-fit events number of dark matter, 8B solar neutrino, AC background, cathode background and MD background in paired and US2 channels, assuming 4 GeV/ DM-nucleon spin-independent interaction and 2 GeV/ LDM-electron interaction via heavy mediator and light mediator.
@@
In the paired dataset, the dominant background arises from accidental pileup (AC) events, whereas the US2 dataset is primarily affected by three components: the cathode-originated background, the pileup of delayed electrons (MD background) 111The acronym “MD” stands for micro-discharging since this kind of background was first observed during a dataset when discharging occurs quite frequently [PandaX:2022xqx].
, and the coherent elastic neutrino-nucleus scattering induced by solar 8B neutrinos (8B CENS).
@@
The models have been tested with a sideband before unblinding as in Ref. [PandaX:2022xqx, PandaX:2022aac].
Systematic uncertainties are assigned to the selection acceptance, BDT efficiency, MD background rate, cathode background rate, and signal response model, following the same methodology as Ref. [bo2024first].
@@
The background models used in this work can be found in the Supplementary Material [SupplementaryMaterial].
The signal response models for LDM and 8B CENS interactions in both the paired and US2 data are constructed using the PandaX-4T signal response framework (bottom panels in Fig. 1) [luo2024signal].
@@
For comparison, we also interpret the results using the charge yields from the Noble Element Simulation Technique (NEST) model [szydagis2022review], leading to slightly more aggressive results.
Both the NR and ER signal model uncertainties are considered in the likelihood analysis with the same treatment in Refs. [bo2024first, PandaX:2022xqx].
In the paired (US2) dataset, we observe one (158) and two (174) events in run0 and run1 [PandaXUS2Data], respectively.
@@
The joint (, ) distributions of the MD and cathode backgrounds are derived from controlled samples in the experimental data, while those for DM and 8B CENS signals are generated by the signal response model, assuming a uniform spatial distribution within the fiducial volume(FV).
Fig. 2 displays the 1 confidence contours for both background components and DM signals with masses of 4 GeV/ (SI DM-n) and 2 GeV/ (LDM-e), alongside the observed US2 event distributions from run0 and run1.
@@
The 90% CL exclusion limit(blue line) with 1 sensitivity(green band) on spin-independent DM-nucleon cross section(left panel), neutron-only(middle panel) and proton-only(right panel) spin-dependent DM-nucleon cross section, together with results from other work [PandaX:2022xqx, PandaX:2022aac, XENON:2024hup, XENON:2019gfn, DarkSide:2018kuk, LUX:2016ggv, PandaX-II:2018woa, PICO:2019vsc].
@@
The 90% CL exclusion limit(blue line) with 1 sensitivity(green band) on DM-electron cross section via heavy(top panel) and light mediator(bottom panel) using constant-W model. The aggressive results using NEST light and charge yields are also overlaid(blue dashed line), as well as results of other experiments [PandaX:2022xqx, XENON:2019gfn, DarkSide:2022knj, adari2025first, aggarwal2025probing], and cosmological predictions from DM vector-portal freeze-in and freeze-out mechanisms [Essig:2015cda, bhattiprolu2024dark].
@@
Figure 3 presents the 90% confidence level (CL) exclusion limits on the DM-nucleon scattering cross sections.
The results establish the most stringent limits for SI, SD (neutron-only), and SD (proton-only) DM-nucleon scatterings in the mass range of approximately [2.5, 5.0], [2.0, 4.6], and [2.0, 3.6] GeV/c2, respectively.
The corresponding exclusion limits at 3 GeV/ are , , and cm2.
@@
The SI limits exhibit stronger upward fluctuations above 3.5 GeV/ because the harder energy spectrum populates the region (5-8 electrons) where we observe a slight event excess.
Below 3.5 GeV/, SI limits show no significant upward fluctuations as the spectra become dominated by MD background.
This trend also holds for the SD limits.
The LDM-electron cross section exclusion limits are improved by approximately 1.5 times for the heavy-mediator case, and 9.3 times for the light-mediator case, in mass range of [50, 10000] MeV/.
The improved sensitivity for light mediators stems from a refined signal model.
Unlike the previous approach [PandaX:2022xqx], which conservatively floored the mean number of photons and electrons to an integer, this work treats it as a continuous value, providing a more accurate spectrum near the ROI threshold critical for LDM–electron interactions mediated by a light mediator.
The limits for LDM with heavy mediators exhibit stronger upward fluctuations than those with light mediators, as heavy-mediator spectra predominantly populate the excess region (5-8 electrons), while light-mediator spectra primarily occupy the MD region (4-5 electrons).
In addition, due to the expected events gathering in the low-energy region where MD is dominant, greatly reduced MD uncertainty(from 130% [PandaX:2022xqx] to 13% for run0 and 16% for run1) also improves the limits by around 2 times in light-mediator case.
The limit now has excluded portion of the cosmologically allowed parameter space from the cosmic microwave background measurements [Essig:2015cda, bhattiprolu2024dark], in the mass range from about 50 MeV/ to 10 GeV/, which is shown in Fig. 4.
@@
Our analysis establishes new exclusion limits at 90% CL for: spin-independent DM-nucleon scattering in the mass range [2.5, 5.0] GeV/, spin-dependent DM-nucleon scattering ([2.0, 4.6] GeV/ for neutron-only coupling, and [2.0, 3.6] GeV/ for proton-only coupling), and LDM-electron scattering (from 100 MeV/ to 10 GeV/).
@@
For the LDM model, our results constrain previously unexplored parameter space given by the cosmological predictions [Essig:2015cda, bhattiprolu2024dark].
@@
—This project is supported in part by grants from National Key R&D Program of China (Nos. 2023YFA1606200, 2023YFA1606201), National Science Foundation of China (Nos. 12090060, 12090061, 12090063, 12325505, 12222505, 12275267, U23B2070), and by Office of Science and Technology, Shanghai Municipal Government (grant Nos. 22JC1410100, 21TQ1400218), Sichuan Province Innovative Talent Funding Project for Postdoctoral Fellows (No. BX202322), Sichuan Provincial Natural Science Foundation (No. 2024NSFSC1374). We thank for the support by the Fundamental Research Funds for the Central Universities. We also thank the sponsorship from the Chinese Academy of Sciences Center for Excellence in Particle Physics (CCEPP), Thomas and Linda Lau Family Foundation in Hong Kong, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund. Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.
—The data that support the findings of this article are openly available[PandaXUS2Data].